OpenStudy (anonymous):

3y-4=2 I have to use the undo thing someone please help.!

OpenStudy (anonymous):

3y-4=2 add 4 to both sides 3y-4+4=2+4 3y=6 Please divide both sides by 3 3y/3=6/3 y=2......Answer

OpenStudy (anonymous):

Y = 2

OpenStudy (anonymous):

ohhh thank you ! :)